OpenStudy (anonymous):

Please help me "Solve square root of the quantity x minus three equals square root of two minus square root of x for x" A:25/8 B:no solution C:-25/8 D:1/8

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \sqrt{x-3}= \sqrt{2} - \sqrt{x} }\) like this ?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \sqrt{x-3}= \sqrt{2} - \sqrt{x} }\) square both sides, and you get \(\normalsize\color{blue}{ x-3= 2 - \sqrt{2x} +x }\) \(\normalsize\color{blue}{ x-3\color{red}{ -2-x }= 2 - \sqrt{2x} +x \color{red}{ -2-x } }\) \(\normalsize\color{blue}{ -5= - \sqrt{2x} }\) \(\normalsize\color{blue}{ 25= 2x }\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ x=? }\)

OpenStudy (anonymous):

The equation was suppose to solve for x, no x value was given.

OpenStudy (anonymous):

does that mean the answer is B, no solution?